### v2

# Sum-product phenomena for Ahlfors-regular sets

Abstract.
We utilise the recent work of Orponen to yield a result of sum and product for Ahlfors-regular sets. The result is sharp up to constants. As a corollary, we obtain the fractal analogue of Solymosi’s -bound for finite subsets of
Key words and phrases:
Discretised sum-product, discretised ring theorem, Shannon entropy, Ahlfors-regular sets2010 Mathematics Subject Classification:
05B99, 28A78, 28A801. Introduction
Sum-product phenomena is the maxim that additive and multiplicative structure find it hard to co-exist. An example of this is the result of Erdős and Szemerédi [ES83]: There exists so that for all large enough we have
In other words, finite subsets of cannot closely resemble rings. The above is now known the hold over and the best known value of one can take is due to Rudnev and Stevens [RS22]; It is conjectured that one may take any in other words, either the sum-set or the product-set must be almost as large as possible.
Sum-product phenomena is now known to hold in a large variety of settings, in particular, for fractal sets. Loosely stated, if has ‘dimension’ , then one of or has ‘dimension’ ‘much larger’ than This problem was introduced by Katz and Tao [KT01] as it is related to the now solved Erdős–Volkmann ring problem [EV66]: are there Borel subrings of with Hausdorff dimension strictly between and This was answered in the negative by Bourgain [Bou03] and Edgar–Miller [EM03] independently.
The Edgar–Miller paper gave a very direct and fairly elementary proof. Bourgain showed that the ring problem is a corollary of the so called ‘discretised ring theorem’: For all there exists so that if is a finite -separated set that resembles a fractal set 111 for all of dimension then
provided that is small enough. This problem has attracted a large amount of interest in recent years. For the best bound on see [FR24], [RW23]. For an elementary proof see [GKZ21]. For a result with weaker conditions on see [BG08], [Bou10].
The aim of this note is to improve on the bound given in [RW23], but for the more restricted class of Ahlfors regular sets. The required definitions will be given in section 2.
Theorem 1.3.
Let and let There exists a so that the following holds. Let be an Ahlfors -regular measure with and let be i.i.d. random variables distributed by We have
for all
Here, and throughout, denotes the Shannon entropy with respect to a partition of intervals of length This will be defined and made precise in section 2. Theorem 1.3 is sharp; see section 4. A corollary is the following.
Theorem 1.5.
Let and let There exists a so that the following holds. Let be an Ahlfors -regular compact set. We have
and
for all
Here denotes the least number of closed intervals of length needed to cover the set. Again, (1.6) is sharp as we will see in section 4. Inequality (1.7) resembles the bound obtained by Solymosi in [Sol09] for the discrete sum-product problem.
1.1. Proof sketch
The remarkable recent result of Orponen shows that for an Ahlfors -regular measure and a -Frostman measure both supported on we may always find so that
As an artifact of his proof, we are able to massage and manipulate his work into the following result. Let be Ahlfors -regular. Let be i.i.d. random variables distributed by Then
We may then use the submodularity of Shannon entropy and the fact that are i.i.d. to show that
Combining these inequalities gives the required result. The use of the submodularity of entropy has been previously utlised in the work of Máthé and the author to gain a strong bound for the discretised ring theorem. See [MO23].
Acknowledgements
I would like to thank Pablo Shmerkin and Joshua Zahl for invaluable discussions and suggestions. I would also like to thank Tuomas Orponen for pointing out [MS09].
2. Preliminaries
We recall what we need.
2.1. Ahlfors and upper Ahlfors regularity
Definition 2.1.
Let A Borel probability measure on is called Ahlfors -regular if
for all and Certainly this means that for all we get
A closed set is Ahlfors -regular if there exists a Borel probability measure which is Ahlfors -regular with
Definition 2.4.
Let A set is called upper -regular if
for all and A Borel probability measure is -Frostman if for all and Further, it is -regular if is upper -regular.
2.2. Finding thick Ahlfors regular subsets
We need a result of Mattila and Saaranen regarding finding thick Ahlfors -regular subsets of Ahlfors -regular sets. [MS09].
Theorem 2.6.
[MS09, Theorem 5.1] Let and Then there exists a constant so that the following holds. Let be Ahlfors -regular. Then there exists a Borel measure with so that
for all and Further,
We will apply it in the following form.
Lemma 2.8.
Consider the same hypothesis as the result above. Then there exists so that the following holds. Let be Ahlfors -regular. Then there exists an Ahlfors -regular measure with
We make the simple deduction.
Proof.
Use the previous theorem to find and the described measure Set If is a probability measure, then we are done. If it is not, we renormalise: Let Using (2.7) we see that
for all and Let Since we have
Adding this to the previous set of inequalities gives us
Now recall that Also, Adding this into the above gives
Taking gives us the required result. ∎
2.3. Entropy
In the below our random variables may take finitely many values in Let be a random variable. Define the Shannon entropy of by
Define the collision entropy of by
Here, and throughout, will be taken to base 2, but this is not particularly important. We adhere to the convention that
We recall some useful facts. The first is monotonicity: For any random variable we have
The second is that Shannon and collision entropy are concave. To see this for collision entropy, note that the -norm is convex, its composition with the concave function leaves it convex, and then taking the negative turns it concave. The below well-known inequality is useful.
Theorem 2.16 (Submodular inequality).
Let be random variables. Suppose that determines and determines Suppose that determines Then
We wish to not restrict ourselves to just random variables which have finite support. But, we do want to use the theory above. We do this by discretising our (infinitely supported) random variables at a scale To this end, let be the collection of -cubes of the form Now let be a compactly supported random variable on Write the -Shannon entropy of by,
and the -collision entropy of by,
We still have and inherit the concavity from the finite setting. We also need the following facts.
Lemma 2.20 (Restriction).
Fix Let be a random variable and suppose that is an event so that Then
Here is the random variable conditioned on
Proof.
Let be the distribution of We may write
By the concavity of collision entropy we then obtain
A lower bound of the left-hand side is
and so we are done. ∎
Let and We say that a random variable is -Frostman if
for all
Lemma 2.25.
Suppose that is -Frostman. Then
Proof.
We have
We desire a submodular inequality in this setting too.
Lemma 2.30 (Discretised submodular inequality).
Let be random variables taking values in compact subsets of respectively. Fix Suppose each of the following:
If we know that the outcome of lies in then we are able to determine which the outcome of will lie;
If we know that the outcome of lies in then we are able to determine which the outcome of will lie;
If we know the outcome of lies in and the outcome of lies in then we are able to determine which the outcome of will lie.
Then,
where the implicit constant depends on only.
Proof.
Define the discrete random variables on the sample space which output the which the outputs of lie in, respectively. Similarly, define the discrete random variables on the sample space which output the which the outputs of lie in, respectively. It is clear that
and
By construction, determines as does and determines Therefore by submodularity with we have
Using the above identifications gives us,
Finally by the continuity of entropy we have the result required. ∎
A useful rendition of this is the following.
Lemma 2.31.
Let be i.i.d random variables taking values in We have
for all Here is a generic constant (it depends on , the support of the random variables).
Proof.
By submodularity, we have
Since are i.i.d. the result follows. ∎
2.4. High multiplicity sets and a result of Orponen
We recall some definitions from Section 2.1 in [Orp24].
Definition 2.34.
[Orp24, Notation 2.1] Let and Define the multiplicity function by
Here is the orthogonal projection of to the line spanned by Also write
We have the recent and remarkable result of Orponen [Orp24].
Theorem 2.37.
[Orp24, Theorem 1.13] For every and there exists so that the following holds. Let be a -regular on and let be -Frostman on Then,
for all
Note that the radius of the ball being is arbitrary, and we may replace with (say), whilst altering our outputted by a factor of at most 10.
This is good, but not exactly what we want. We would like a result where orthogonal projections are replaced with radial projections with centres contained on a line. This is possible via a projective transformation. We restate Definition 2.34 in this setting.
Definition 2.39.
Let and Define the radial multiplicity function by
Here is the radial projection to the circle of radius 1 centred at Also write
Lemma 2.42.
For every and there exists so that the following holds. Let be a -regular on and let be -Frostman with where is a line. Suppose further that Then,
for all
We only apply this lemma when so we only complete the proof in this case. The general case can be easily proved by modifying what follows. See also [OSW24, Remark 4.13].
Proof of Lemma 2.42.
Define the projective transformation by
For and let The family
contains all the lines passing through which are not contained in It is easy to see that where Therefore, transforms lines in to lines parallel to the vector Now extend to map a point on the line to the line at infinity spanned by
We wish to apply Theorem 2.37 to the transformed measures and We just need to check that these measures are still regular with reasonable constant. This will follow from the fact that is bi-Lipschitz when restricted to which, in turn, is due to the separation of the measures and and the support of being contained in Let be the bi-Lipschitz constant of (set for example). We have, by applying
where Applying Theorem 2.37 gives us that the right-hand side, and therefore the left-hand side, is for all and therefore small enough, that depends only on as required. ∎
We prove an entropic version of the above.
Proposition 2.46.
For every there exists so that the following holds. Let be -regular on and let be -Frostman on Suppose that Let be independent random variables distributed by respectively. Then
Proof.
By the definition of conditional entropy we have
Fix We examine By monotonicity of Renyi entropy we know that
Write
Let be a trial distributed by
By the restriction estimate we have
We now lower-bound Let be the tubes coming from the pull-backs of a -covering of the radial projection Since is -regular we certainly have that Consider such a tube Take a line contained in Since can be covered by balls of radius the tube can be covered by such balls. Each ball has measure at most Putting these two facts together tells us that the tube has measure at most We now estimate the collision entropy.
Therefore
Combining with the restriction estimate above and monotonicity of Renyi entropy we have
This gives us
as required. ∎
Corollary 2.56.
Let be Ahlfors -regular on . Let be i.i.d. random variables distributed by We have
for all
Proof.
Apply Proposition 2.46 with the measures and (The map is bi-Lipschitz when restricted to so it is the case that is Ahlfors -regular, where we may need to increase by a factor of ) ∎
3. Proof of Theorem 1.3
Proof of Theorem 1.3.
Fix If then choose so small so that
Applying Corollary 2.56 with gives us
If then we use Lemma 2.8 to find a constant and an Ahlfors -regular measure with Let be i.i.d. random variables distributed by Again, applying Corollary 2.56 with we obtain
Now take smaller until becomes and we are done. ∎
4. Sharpness of Theorem 1.3
We use the language of iterated function systems, see chapter 11 of [Fal14]. Fix Let be an arithmetic progression of length starting from and of step Consider the iterated function system where is chosen so that Let be the attractor of and let be the self-similar measure with uniform weights on It is well known that is Ahlfors -regular for some Further, for all small enough, and large enough, we have Now will depend on and Consider the map This restricted to is bi-Lipschitz and its image is contained in The image measure of is therefore still Ahlfors -regular, where depends on only. Also, for we have that for all small enough. Let be i.i.d. random variables distributed by For such a we have