### v3

# Sum-product phenomena for Ahlfors-regular sets

Abstract.
We utilise the recent work of Orponen to yield a sum-product result for Ahlfors-regular sets. As a corollary, we obtain the fractal analogue of Solymosi’s -bound for finite subsets of
Key words and phrases:
Discretised sum-product, discretised ring theorem, Shannon entropy, Ahlfors-regular sets2010 Mathematics Subject Classification:
05B99, 28A78, 28A801. Introduction
The sum-product phenomenon is the maxim that additive and multiplicative structure find it hard to co-exist. An example of this is the result of Erdős and Szemerédi [ES83]: There exists so that for all large enough we have
In other words, finite subsets of cannot closely resemble rings. The above is now known to hold over and the best known value of (in print) one can take is due to Rudnev and Stevens [RS22]; (See also the recent preprint [Blo25] that gives a modest improvement to the work of Rudnev–Stevens). It is conjectured that one may take any in other words, either the sum-set or the product-set must be almost as large as possible.
Sum-product phenomena are now known to hold in a large variety of settings, in particular, for fractal sets. Loosely stated, if has ‘dimension’ , then one of or has ‘dimension’ ‘much larger’ than This problem was introduced by Katz and Tao [KT01] as it is related to the now solved Erdős–Volkmann ring problem [EV66]: are there Borel subrings of with Hausdorff dimension strictly between and This was answered in the negative by Bourgain [Bou03] and Edgar–Miller [EM03] independently.
The Edgar–Miller paper gave a very direct and fairly elementary proof. Bourgain showed that the ring problem is a corollary of the so called ‘discretised ring theorem’: For all there exists so that if is a finite -separated set that resembles a fractal set 111 for all of dimension then
provided that is small enough. Here, and throughout, denotes the least number of closed intervals of length needed to cover the set. This problem has attracted a large amount of interest in recent years. For the best bound on see [FR24], [RW23]. For an elementary proof see [GKZ21]. For a result with weaker conditions on see [BG08], [Bou10].
The aim of this note is to improve on the bound given in [RW23], but for the more restricted class of regular sets. See the below definition.
Definition 1.3.
Let A set is called upper -regular if
for all and
A Borel probability measure is -Frostman if for all and Further, it is -regular if is upper -regular.
A set is -regular if there exists an -regular measure with
Theorem 1.5.
Let and let There exists a so that the following holds. Let be an -regular measure with and let be i.i.d. random variables distributed by We have
for all
Here, and throughout, denotes the Shannon entropy with respect to a partition of intervals of length This will be defined and made precise in section 2. Theorem 1.5 is sharp in the sense that cannot be replaced by a larger number; see section 4. We remark that in Theorem 1.5, we may replace with and/or replace with by minor alterations in the proof. An immediate corollary follows using the fact that
Theorem 1.7.
Let and let There exists a so that the following holds. Let be an -regular set. We have
and
for all
Again, (1.8) is sharp as we will see in section 4, and we may replace with and/ or replace with Inequality (1.9) resembles the bound obtained by Solymosi in [Sol09] for the discrete sum-product problem, although the proofs seem totally unrelated.
1.1. Proof sketch
Fix Here we use loosely to morally mean greater than. A remarkable recent result of Orponen [Orp24] shows that for an -regular measure and a -Frostman measure both supported on we may always find so that
As an artifact of his proof, we are able to massage and manipulate his work into the following result. Let be Ahlfors -regular. Let be i.i.d. random variables distributed by Then
We may then use the submodularity of Shannon entropy and the fact that are i.i.d. to show that
Combining these inequalities gives the required result. The use of the submodularity of entropy has been previously utlised in the work of Máthé–O’Regan to gain a strong bound for the discretised ring theorem. See [MO23].
Acknowledgements
We thank Pablo Shmerkin and Joshua Zahl for invaluable discussions and suggestions, in particular to Joshua for pointing out that Theorem 1.5 is sharp. We thank Tuomas Orponen for useful comments.
2. Preliminaries
We recall what we need.
2.1. Entropy
In the below our random variables may take finitely many values in Let be a random variable. Define the Shannon entropy of by
Define the collision entropy of by
Here, and throughout, will be taken to base 2, but this is not particularly important. We adhere to the convention that
We recall some useful facts. The first is monotonicity: For any random variable we have
The second is that Shannon and collision entropy are concave. To see this for collision entropy, note that the -norm is convex, its composition with the concave function leaves it convex, and then taking the negative turns it concave. The third is the chain rule. Let be random variables, then
The below well-known inequality is useful. For example, see [Tao10, Lemma A.2]. For two random variables we say that determines if when the outcome of is known, then so is the outcome of
Theorem 2.5 (Submodular inequality).
Let be random variables. Suppose that determines and determines Suppose that determines Then
We wish to not restrict ourselves to just random variables which have finite support. But, we do want to use the theory above. We do this by discretising our (infinitely supported) random variables at a scale To this end, let be the collection of -cubes of the form For we let denote the intervals of which intersect Now let be a compactly supported random variable on Write the -Shannon entropy of by,
and the -collision entropy of by,
We still have
and inherit the concavity from the finite setting. We also need the following facts.
Lemma 2.10 (Continuity).
Suppose that is a random variable on a compact set and let , Then
with the implicit constant depending on only.
Proof.
Let be the random variable on which outputs the for which let be the random variable on which outputs the for which We have by the chain rule (2.4),
which leads us to
Finally, for each the random variable has a sample space of size at most and so for each we have
and so taking expectation gives us
and the result follows. ∎
Lemma 2.11 (Restriction).
Fix Let be a random variable and suppose that is an event with Then
Here is the random variable conditioned on
Proof.
Let be the distribution of We may write
By the concavity of collision entropy we then obtain
A lower bound of the left-hand side is
and so we are done. ∎
Let and We say that a random variable is -Frostman if
for all
Lemma 2.16.
Suppose that is -Frostman. Then
Proof.
We have
We desire a submodular inequality in this setting too, which has previously been obtained in [MO23]. We include their proof for convenience.
Lemma 2.21 (Discretised submodular inequality).
Let be random variables taking values in compact subsets of respectively. Fix Suppose each of the following:
If we know that the outcome of lies in then we are able to determine a choice of which the outcome of will lie;
If we know that the outcome of lies in then we are able to determine a choice of which the outcome of will lie;
If we know the outcome of lies in and the outcome of lies in then we are able to determine a choice of which the outcome of will lie.
Then,
where the implicit constant depends on only.
Proof.
Define the discrete random variables on the sample space which output the which the outputs of lie in, respectively. Similarly, define the discrete random variables on the sample space , respectively which output the which the outputs of lie in, respectively. It is clear that
and
By construction, determines as does and determines Therefore by submodularity (Theorem 2.5) with we have
Using the above identifications gives us,
Finally by the continuity of entropy (Lemma 2.10) we have the result required. ∎
A useful rendition of this is the following.
Lemma 2.22.
Let be i.i.d random variables taking values in We have
for all
Proof.
We essentially use the fact that the maps are Lipschitz with constant (say), when restricted to and the map is Lipschitz with constant when restricted to
Using the facts above, suppose we know a ball of diameter in which lies. Then we are able to locate an interval of length in which will lie. Similarly, suppose we know a ball of diameter in which lies, then we are able to locate an interval of length in which will lie.
Now suppose we jointly know two balls of diameter where lies in one, and lies in the other. We therefore can determine an interval of length in which lies, and so two intervals of length one containing and containing
3. Proof of Theorem 1.5
3.1. High multiplicity sets and a result of Orponen
We recall some definitions from [Orp24, Section 2.1].
Definition 3.1.
[Orp24, Notation 2.1] Let and Define the multiplicity function by
Here is the -neighbourhood of and is the orthogonal projection of to the line spanned by Also write
We have the recent and remarkable result of Orponen [Orp24].
Theorem 3.4.
[Orp24, Theorem 1.13] For every and there exists so that the following holds. Let be a -regular measure on and let be -Frostman on Then,
for all
This is good, but not exactly what we want. We would like a result where orthogonal projections are replaced with radial projections with centres contained on a line. This is possible via a projective transformation. We restate Definition 3.1 in this setting.
Definition 3.6.
Let and Define the radial multiplicity function by
Here is the radial projection to the circle of radius 1 centred at Also write
Lemma 3.9.
For every and there exists so that the following holds. Let be an -regular on measure and let be an -Frostman measure with where is a line. Suppose further that and Then,
for all
We only apply this lemma when so we only complete the proof in this case. The general case can be easily proved by modifying what follows. See also [OSW24, Remark 4.13].
Proof of Lemma 3.9.
Define the projective transformation by
For and let The family
contains all the lines passing through which are not contained in It is easy to see that where Therefore, transforms lines in to lines parallel to the vector Now extend to map a point on the line to the line at infinity spanned by
We wish to apply Theorem 3.4 to the transformed measures and We just need to check that these measures are still regular with reasonable constant. This will follow from the fact that is bi-Lipschitz when restricted to which, in turn, is due to the separation of the measures and and the support of being contained in Let be the bi-Lipschitz constant of (set for example). We have, by applying
where Applying Theorem 3.4 gives us that the right-hand side, and therefore the left-hand side, is for all and therefore small enough, that depends only on as required. ∎
We prove an entropic version of this below. Fix We wish to identify the radial projection with the map which measures the slope between and i.e This comparison will make sense provided that bounded away from The fact we will need is the following: Let Let be an interval of length in Then
provided is small enough.
Proposition 3.14.
For every there exists so that the following holds. Let be -regular on and let be -Frostman on Suppose that and
Let be independent random variables distributed by respectively. Then
Proof.
Let be that given by Lemma 3.9. Fix By the definition of conditional entropy we have
Fix We examine By monotonicity (2.9) we know that
Write
Let be a trial distributed by
By Lemma 2.11 we have
We now lower-bound Let be the tubes coming from the pull-backs of a -covering of Note that by (3.13) these tubes will have width Since is -regular we certainly have that Consider such a tube Take a line contained in Since can be covered by balls of radius the tube can be covered by such balls. Each ball has measure Putting these two facts together tells us that the tube has measure at We now estimate the collision entropy
Therefore
Combining with the restriction estimate (Lemma 2.11) and monotonicity (2.9) we have
By Lemma 3.9, we have
and using the fact that we have
This gives us
as required. ∎
Corollary 3.27.
Let be -regular on . Let be i.i.d. random variables distributed by We have
for all
Proof.
Apply Proposition 3.14 with the measures and (The map is bi-Lipschitz when restricted to so it is the case that is -regular, where we may need to increase by a factor of ) ∎
3.2. Proof of Theorem 1.5
4. Sharpness of Theorem 1.5
We use the language of iterated function systems, see [Fal14, Chapter 11]. Fix Let be an arithmetic progression of length starting from and of step Consider the iterated function system where is chosen so that Let be the attractor of and let be the self-similar measure with uniform weights on It is well known that is -regular for some Further, for all small enough, and large enough, we have Now will depend on and Consider the map This restricted to is bi-Lipschitz and its image is contained in The image measure of under the map is therefore still -regular, where depends on only. Also, for we have that for all small enough. Let be i.i.d. random variables distributed by For such a we have